OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 7 cm. What is the approximate radius of the cylinder? A. 4.0 cm B. 3.5 cm C. 2.2 cm D. 1.5 cm HELP

OpenStudy (gudden):

volume ofcylinder = pi r^2 h where h = height of the cylinder r = radius of the cylinder and volume of rectangular prism = l.b.h. where l = length b = breadth and h = height and since as it has been given that its a square with side length 7cm therefore l *b= 7*7 now try equating these two equations

OpenStudy (gudden):

if u still find a problem.. just let me know

OpenStudy (anonymous):

I got b